A company just paid a $2 dividend per share. The dividend growth rate is expected to be constant at 10% for 3 years, after which dividends are expected to grow at a rate of 4% forever. If the company’s required return (rs) is 10%, what is its current stock price?
Answer #1

D1=(2*1.1)=2.2

D2=(2.2*1.1)=2.42

D3=(2.42*1.1)=2.662

Value after year 3=(D3*Growth Rate)/(Required rate-Growth Rate)

=(2.662*1.04)/(0.1-0.04)

=$46.1413(Approx)

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.2/1.1+2.42/1.1^2+2.662/1.1^3+$46.1413/1.1^3

=$40.67(Approx).